OpenStudy (anonymous):

Which of the following graphs would help find the solution to this system equations? x + y = -4 4x + y = -18 answer choices will be listed in the post

OpenStudy (anonymous):

info

OpenStudy (goten77):

x+y=-4 same as y=-x-4 4x+y=-18 same as y=-4x-18 now set each right side equal to eachother -x-4 = -4x-18 +x +x -4 = -3x -18 +18 +18 14= -3x /-3 /-3 x=14/-3 =-4.6 now we know that they intercept at a point where x=-4.6 which is only on 1 of the graphs :s